OpenStudy (anonymous):

Provide the solution to the system using elimination. 2x – 5y = –12 3x – 4y = 3 Please show me the work and the answer

OpenStudy (anonymous):

we can do this together :)

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

first multiply 1st equation with 3 and 2nd equation with 2

OpenStudy (anonymous):

wait but there are only two equations

OpenStudy (anonymous):

Yes. Multiply the first equation by 3. Multiply the second equation by 2. Then the x variable will have the same coefficient in each.

OpenStudy (anonymous):

And you can subtract one from the other and eliminate the x's

OpenStudy (anonymous):

whats 3(2x-5y)=3(-12)

OpenStudy (anonymous):

-36?

OpenStudy (anonymous):

do you know whats 3(2x-5y)

OpenStudy (anonymous):

can you ?

OpenStudy (anonymous):

6x-15y

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

so,6x-15y=-36

OpenStudy (anonymous):

2(3x-4y)=2(3)

OpenStudy (anonymous):

isnt it 6x-8y=6?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

1st equation is 6x-15y=-36 2nd equation is 6x-8y=6

OpenStudy (anonymous):

now subtract 2nd from 1st

OpenStudy (anonymous):

6x-15y=-36 - 6x-8y=6

OpenStudy (anonymous):

so, 6x-6x-15y+8y=-36-6

OpenStudy (anonymous):

is the answer 6?

OpenStudy (anonymous):

yea y=6 now just plug y value in one of the equation to get x value

OpenStudy (anonymous):

which equation the first one?

OpenStudy (anonymous):

as you like

OpenStudy (anonymous):

2x-5(6)=-12

OpenStudy (anonymous):

3x+4(6)=3?

OpenStudy (anonymous):

no,3x-4(6)=3

OpenStudy (anonymous):

oh so then I solve ti like a normal equation right?

OpenStudy (anonymous):

it*

OpenStudy (anonymous):

Now you're solving for x. Once you have your x and your y, you put them together to make the solution (x,y)

OpenStudy (anonymous):

so x=-7 and y=6 right?

OpenStudy (anonymous):

nope, 3x-4(6)=3 3x-24=3 3x=27 x=27/3 x=9

OpenStudy (anonymous):

oh ok thanks alot :)